# One Demo Is All It Takes: Planning Domain Derivation with LLMs from A Single Demonstration

### v3

@@ Abstract
Pre-trained large language models (LLMs) show promise for robotic task planning but often struggle to guarantee correctness in long-horizon problems. Task and motion planning (TAMP) addresses this by grounding symbolic plans in low-level execution, yet it relies heavily on manually engineered planning domains. To improve long-horizon planning reliability and reduce human intervention, we present Planning Domain Derivation with LLMs (PDDLLM), a framework that automatically induces symbolic predicates and actions directly from demonstration trajectories by combining LLM reasoning with physical simulation roll-outs. Unlike prior domain-inference methods that rely on partially predefined or language descriptions of planning domains, PDDLLM constructs domains without manual domain initialization and automatically integrates them with motion planners to produce executable plans, enhancing long-horizon planning automation. Across 1,200 tasks in nine environments, PDDLLM outperforms six LLM-based planning baselines, achieving at least 20% higher success rates, reduced token costs, and successful deployment on multiple physical robot platforms.
@@ 1 Introduction
Robotic planning remains challenging in complex scenarios that require abstract, long-horizon reasoning. Large language models (LLMs) demonstrate strong generalization in this domain but often struggle with temporal dependencies in extended tasks Huang et al. (2022, 2023). Task and Motion Planning (TAMP) frameworks provide robustness in long-horizon reasoning by integrating high-level symbolic reasoning with low-level motion planning. However, they face two major limitations: (i) planning domains, expressed in symbolic planning languages such as PDDL (McDermott et al., 1998), are difficult to ground in complex geometric information, and (ii) these domains are labor-intensive to manually construct Garrett et al. (2020, 2021); Khodeir et al. (2023); Silver et al. (2020). Recent vision-language-action models have advanced the first challenge by improving geometric grounding through semantic action instructions, but they pay limited attention to long-horizon reasoning Black et al. (2024); Kim et al. (2024); Zitkovich (2023); Li et al. (2024); Team (2024). Complementarily, our work focuses on solving the second challenge by reducing the human effort required for domain construction, thereby enabling automated and scalable long-horizon reasoning.
Existing domain-generation methods typically rely on predesigned elements (predicates or actions) to complete a domain Silver et al. (2023); Kumar et al. (2023), where manual predesign remains critical and often depends on curated training data Huang et al. (2025). LLM-based approaches also demand extensive natural-language descriptions of PDDL domains and careful prompt engineering Guan et al. (2023a); Oswald et al. (2025). Meanwhile, recent advances in world-model learning show that LLMs can serve as compact explainers and reasoners for robot behavior and object relations (Guan et al., 2023b; Zhao et al., 2024; Liang et al., 2024; Tang et al., 2024). Motivated by this, we propose to combine LLM reasoning with physical simulation to generate planning domains, eliminating reliance on manual predesign, curated training data, and detailed textual descriptions.
We introduce PDDLLM, the first method to construct planning domains without manual predesign, training data, or extensive human descriptions. Moreover, PDDLLM automates integration with low-level motion planners, further reducing the need for manual intervention. Our method targets at long-horizon planning problems, aligning with the scope of TAMP studies. Managing action sequences and task structures is essential for these tasks such as household activities and puzzle solving. The main contributions of this paper are:
@@ 2 Related Work
With the advent of pre-trained large models (LMs), the use of LLMs and vision language models (VLMs) has significantly advanced the performance of task planners (Huang et al., 2023; Wang et al., 2024a; Chen et al., 2024; Li et al., 2023). Although many studies have demonstrated the generalization capabilities of LM-based task planners, they often lack robustness and struggle with long-horizon tasks that require complex reasoning (Wang et al., 2024a; Sermanet et al., 2023; Driess et al., 2023; Brohan et al., 2023; Huang et al., 2022). To address this limitation, recent research has explored guiding task planners with LM-derived heuristics to accelerate informed search. These approaches integrate symbolic search with LMs to accelerate task planning and reduce search complexity. Notable efforts include heuristics for prioritizing feasible states (Zhao et al., 2024), ranking feasible actions (Yang et al., 2024; Meng et al., 2024; Hu et al., 2023; Zhao et al., 2024), and pruning search trees (Silver et al., 2024). However, a major limitation of these methods lies in their reliance on manually constructed symbolic planning domains to build search trees, which imposes additional development overhead and reduces flexibility.
@@
A recent line of research aims to infer the planning domain directly from human demonstrations, environment interactions, or natural language. However, these approaches often depend on partially or fully predefined symbolic predicates and actions as well as extensive training dataset (Silver et al., 2023; Kumar et al., 2023; Huang et al., 2025; Wong et al., 2023; Liu et al., 2025). Some recent studies have explored leveraging LMs for domain generation, primarily by extending manually defined domains with additional predicates and actions (Liang et al., 2024; Athalye et al., 2024; Byrnes et al., 2024). The approaches proposed by Guan et al. (2023b); Han et al. (2024) generates planning domains requiring prompts containing human-crafted planning examples or detailed descriptions of predicates and actions. Moreover, many of these studies assume that symbolic actions must be manually linked to existing motion-level skills (Kumar et al., 2023; Huang et al., 2025; Liang et al., 2024; Athalye et al., 2024), requiring labor-intensive alignment between planning domains and low-level motion planners. LoCA addresses this limitation by automatically grounding symbolic actions into motion constraints, thereby reducing manual effort and enhancing scalability and autonomy.
@@ 3 Preliminaries
PDDL is a standard formal language used to specify planning problems. The object set $\mathcal{O}$ represents the environment’s objects, whose continuous state $\mathcal{S}$, such as pose, color, and size, can be queried via a perception function $\mathcal{I}:\mathcal{O}\times\mathcal{I}\rightarrow\mathcal{S}$. The PDDL domain $\mathcal{D}=(\mathcal{P},\mathcal{A})$ describes the general rules of the environment, consisting of a set of logical predicates $\mathcal{P}$ and a set of logical actions $\mathcal{A}$. A logical predicate $p\in\mathcal{P}$ specifies either intrinsic properties of an object $o$ or relations between objects (e.g.,(cooked ?o1), (on ?o1 ?o2)). Each predicate is evaluated by a binary classifier over the continuous state, returning true or false. A symbolic descripton of the environment $\mathcal{X}$ can be obainted by grounding $\mathcal{P}$ across $\mathcal{S}$ (i.e. $\mathcal{S}\times\mathcal{P}\rightarrow\mathcal{X}$). A logical action $a\in\mathcal{A}$ consists of a precondition $\mathcal{P}_{pre}$ $\langle p_{1},p_{2},\dots\rangle$ and an action’s effect $\mathcal{P}_{eff}=\langle p_{1}^{\prime},p_{2}^{\prime},\dots\rangle$. The precondition represents a set of predicates that must be satisfied for the action to be executed, while the action’s effect describes the change of the resultant state upon action completion. Logical actions define the logical state transitions $\mathcal{X}^{(t)}\times a^{(t)}\rightarrow\mathcal{X}^{(t+1)}$. Thus, any planning problem $\langle\mathcal{S}^{(init)},\mathcal{S}^{(final)}\rangle$ can be formulated as a logical planning problem $\mathcal{Q}=\langle\mathcal{O},\mathcal{D},\mathcal{X}^{(init)},\mathcal{X}^{(final)}\rangle$, which is solved by a symbolic planner to produce a task plan ${a^{(0)},a^{(1)},\dots,a^{(T-1)}}=\texttt{PDDLSolver}(\mathcal{Q})$. Each logical action $a$ must then be integrated with corresponding motion planning skills to generate continuous robotic actions $\tilde{a}$ for execution.
@@ 5 Methodology
Figure 1: Overview of the proposed framework. (1) Human demonstrations, in the form of manipulation trajectories, and the corresponding task descriptions, serve as input. Implementation details is shown in Section˜A.12. (2) PDDLLM initiates thousands of parallel simulations to imagine predicates by summarizing the simulation roll-outs with LLMs, returning the predicate library along with the relevance of each predicate to the current task. (3) Actions are invented by an LLM that summarizes logical state transition patterns from the demonstration, which is grounded into logical states using the imagined predicates. (4) The predicates and actions are compiled into a planning domain, which is automatically interfaced with motion planning algorithm by the Logical Constrain Adapter (LoCA) to solve new tasks.
@@ 5.1 PDDL Domain Generation
Given a human demonstration and the corresponding task description, we combine simulated physical interaction and LLM reasoning to produce an executable PDDL planning domain through predicate imagination and action invention.
@@ 5.1.1 Predicate Imagination
Figure 2: a. This example illustrates the imagination of predicates for relative object positions. Let $u$ be a configurable variable determined by the feature subspace size. Object poses are sampled and simulated, with infeasible cases filtered out. Feasible subspaces are provided to the LLM to generate first-order predicates with their corresponding physical constraints. Higher-order predicates can be further derived using logical operators (e.g., "not", "for all") from first-order predicates. Diverse predicate examples are provided in Section˜A.9 b. This example shows how the Stack action is invented. Continuous states are grounded into logical states using the imagined predicates, where the state transition represents the logical action. By prompting the LLM with the pair of the current state and the next state after the transition, we obtain the PDDL definition of the action Stack. c. The integration of actions with the motion planner is handled automatically by LoCA, which retrieves the physical constraints associated with each first-order predicate in the action effect set $\mathcal{P}_{eff}$ and applies these constraints for motion planning. Prompt examples for a and b are provided in Section˜A.3.
Stage 2. First-order predicates are combined with logical operators and quantifiers to construct higher-order predicates. They are systematically combined in all possible ways to generate more complex logical expressions. Following prior work (Curtis et al., 2022; Silver et al., 2023), we primarily use the negation operator, along with the quantifiers “for all” ($\forall$) and “there exists” ($\exists$). For example, (is_on ?o1, ?o2) can be negated to produce (not_is_on ?o1, ?o2). When combined with the universal quantifier, this further yields ($\mathrm{}{\forall}$_o1_not_is_on ?o1, ?o2), meaning that for any o1 in the environment, (is_on ?o1, ?o2) is false. This indicates that object o2 is on top.
@@ 5.1.2 Action Invention
5.1.3 Generation Refinement
To improve the quality and accuracy of the generated planning domains, we adopt a parallel prompting strategy. Specifically, multiple candidate PDDL models are generated from the same demonstration, and candidates that fail to produce feasible task plans are replaced by alternatives. Empirically, we find that when the number of parallel prompts exceeds five, the correctness of the final domain converges to a stable level. Examples of the parallel prompting procedure are shown in Section˜A.7
@@ 6 Experiments and Baselines
Building on evaluation practices in TAMP, we consider tasks with concrete conditions such as object position, orientation, size, and color (Silver et al., 2023; Garrett et al., 2020; Kumar et al., 2023; Liang et al., 2024; Huang et al., 2025). We structure our experiments along three dimensions, Task Diversity, Task Complexity, and Knowledge Transferability, whose definitions are provided in Section˜A.10.1, with task details described in Section˜A.4. Our experiments are conducted in PyBullet Simulation (Coumans and Bai, 2021), with the symbolic solver from PDDLStream (Garrett et al., 2020) and the motion planning algorithm from PyBullet-Planning (Garrett et al., 2015; Garrett, 2018). The LLM we used is GPT-4o (OpenAI, 2023).
@@ 6.1 Baselines and Ablations
We implement six baselines and one ablation of PDDLLM to comprehensively evaluate our method. GPT-4o is used as the default LLM unless otherwise specified, and the motion planning algorithm is kept the same for all methods. The complete details are provided in Section˜A.2
  • LLMTAMP, o1-TAMP, R1-TAMP: LLM-based task and motion planning (LLMTAMP) inspired by Huang et al. (2022); Li et al. (2022), which use LLMs for task planning, with language description of planning task as input. The task execution demonstration, same as those used in PDDLLM, was provided in the form of natural language in prompt. In addition to GPT-4o, reasoning LLMs, OpenAI’s o1 (OpenAI, 2024) and Deepseek’s R1 (DeepSeek-AI, 2025), are used as backbones for o1-TAMP and R1-TAMP, respectively.
  • LLMTAMP-$F$: Following the method by Huang et al. (2023); Chen et al. (2023), LLMTAMP-$F$ extends LLMTAMP with a failure feedback loop.
  • LLMTAMP-FR: Following Wang et al. (2024b), LLMTAMP-FR extends failure detection by providing specific failure reasons to guide replanning with the LLM.
@@
Robot planning is required to be real-time in robot deployment, imposing constraints on the planning time allowance (Gammell et al., 2015; Garrett et al., 2015, 2020). In our experiment, a uniform planning time limit of 50 seconds is applied to all planning problems and methods. We measure performance using the planning success rate, as in other studies (Huang et al., 2025; Silver et al., 2023; Kumar et al., 2023). The planning time and token cost are used to measure the planning cost (Zhong et al., 2024). Three parallel runs were conducted to compute the mean and standard error for the planning success rate.

### v4

@@ Abstract
Pre-trained large language models (LLMs) show promise for robotic task planning but often struggle to guarantee correctness in long-horizon problems. Task and motion planning (TAMP) addresses this by grounding symbolic plans in low-level execution, yet it relies heavily on manually engineered planning domains. To improve long-horizon planning reliability and reduce human intervention, we present Planning Domain Derivation with LLMs (PDDLLM), a framework that automatically induces symbolic predicates and actions directly from demonstration trajectories by combining LLM reasoning with physical simulation roll-outs. Unlike prior domain-inference methods that rely on partially predefined or language descriptions of planning domains, PDDLLM constructs domains with minimal manual domain initialization and automatically integrates them with motion planners to produce executable plans, enhancing long-horizon planning automation. Across 1,200 tasks in nine environments, PDDLLM outperforms six LLM-based planning baselines, achieving at least 20% higher success rates, reduced token costs, and successful deployment on multiple physical robot platforms.
00footnotetext: Correspondence to: jinbang.huang.work@gmail.com
@@ 1 Introduction
Robotic planning remains challenging in complex scenarios that require abstract, long-horizon reasoning. Large language models (LLMs) demonstrate strong generalization in this domain but often struggle with temporal dependencies in extended tasks (Huang et al., 2022a; b). Task and motion planning (TAMP) frameworks provide robustness in long-horizon reasoning by integrating high-level symbolic reasoning with low-level motion planning. However, they face two major limitations: (i) planning domains, expressed in symbolic planning languages such as PDDL (McDermott et al., 1998), are difficult to ground in complex geometric information, and (ii) these domains are labor-intensive to manually construct (Garrett et al., 2020; 2021; Khodeir et al., 2023; Silver et al., 2021). Recent vision-language-action models have advanced the first challenge by improving geometric grounding through semantic action instructions, but they pay limited attention to long-horizon reasoning (Black et al., 2024; Kim et al., 2024; Zitkovich et al., 2023; Li et al., 2024; Team and others, 2024). Complementarily, our work focuses on solving the second challenge by reducing the human effort required for domain construction, thereby enabling automated and scalable long-horizon reasoning.
Existing domain-generation methods typically rely on predesigned elements (predicates or actions) to complete a domain (Silver et al., 2023; Kumar et al., 2023), where manual predesign remains critical and often depends on curated training data (Huang et al., 2025). LLM-based approaches also demand extensive natural-language descriptions of PDDL domains and careful prompt engineering (Guan et al., 2023; Oswald et al., 2024). Meanwhile, recent advances in world-model learning show that LLMs can serve as compact explainer and reasoner for robot behavior and object relations (Guan et al., 2023; Zhao et al., 2023; Liang et al., 2024; Tang et al., 2024). Motivated by this, we propose to combine LLM reasoning with physical simulation to generate planning domains, eliminating reliance on manual predesign, curated training data, and detailed textual descriptions.
We introduce PDDLLM, an LLM-driven framework for automated planning-domain construction that requires minimal pre-design and no extensive human-provided descriptions. From a single demonstration, the model generate both predicates and actions in a one-shot fashion, yielding an executable planning domain. Moreover, PDDLLM automates integration with low-level motion planners, further reducing the need for manual intervention. Our method targets at long-horizon planning problems, aligning with the scope of TAMP studies. Managing action sequences and task structures is essential for these tasks such as household activities and puzzle solving. The main contributions of this paper are:
@@ 2 Related Work
With the advent of pre-trained large models (LMs), the use of LLMs and vision language models (VLMs) has significantly advanced the performance of task planners (Huang et al., 2022b; Wang et al., 2024; Chen et al., 2024b; Li et al., 2023). Although many studies have demonstrated the generalization capabilities of LM-based task planners, they often lack robustness and struggle with long-horizon tasks that require complex reasoning (Wang et al., 2024; Sermanet et al., 2024; Driess et al., 2023; Ahn et al., 2022; Huang et al., 2022a). To address this limitation, recent research has explored guiding task planners with LM-derived heuristics to accelerate informed search. These approaches integrate symbolic search with LMs to accelerate task planning and reduce search complexity. Notable efforts include heuristics for prioritizing feasible states (Zhao et al., 2023), ranking feasible actions (Yang et al., 2025; Meng et al., 2024; Hu et al., 2023; Zhao et al., 2023), and pruning search trees (Silver et al., 2024). However, a major limitation of these methods lies in their reliance on manually constructed symbolic planning domains to build search trees, which imposes additional development overhead and reduces flexibility.
@@
A recent line of research aims to infer the planning domain directly from human demonstrations, environment interactions, or natural language. However, these approaches often depend on partially or fully predefined symbolic predicates and actions as well as extensive training dataset (Silver et al., 2023; Kumar et al., 2023; Huang et al., 2025; Wong et al., 2023; Liu et al., 2024). Some recent studies have explored leveraging LMs for domain generation, primarily by extending manually defined domains with additional predicates and actions (Liang et al., 2024; Athalye et al., 2024; Byrnes et al., 2025). The approaches proposed by Guan et al. (2023); Han et al. (2024) generates planning domains requiring prompts containing human-crafted planning examples or intense manual feedback.Another line of research investigates generating and refining planning domains through environmental feedback. (Liang et al., 2024; Zhu et al., 2025). Moreover, many studies assume that logical actions have pre-designed motion-level primitive skills (Kumar et al., 2023; Huang et al., 2025; Athalye et al., 2024; Han et al., 2024), requiring labor-intensive alignment between planning domains and low-level motion planners. PDDLLM addresses this limitation by automatically grounding symbolic actions into motion constraints, thereby reducing manual effort and enhancing scalability and autonomy.
@@ 3 Preliminaries
PDDL is a standard formal language used to specify planning problems. The object set $\mathcal{O}$ represents the environment’s objects, whose continuous state $\mathcal{S}$, such as pose, color, and size, can be queried via a perception function $\mathcal{I}:\mathcal{O}\times\mathcal{I}\rightarrow\mathcal{S}$. The PDDL domain $\mathcal{D}=(\mathcal{P},\mathcal{A})$ describes the general rules of the environment, consisting of a set of logical predicates $\mathcal{P}$ and a set of logical actions $\mathcal{A}$. A logical predicate $p\in\mathcal{P}$ specifies either intrinsic properties of an object $o$ or relations between objects (e.g.,(cooked ?o1), (on ?o1 ?o2)). Each predicate is evaluated by a binary classifier over the continuous state, returning true or false. A symbolic description of the environment $\mathcal{X}$ can be obtained by grounding $\mathcal{P}$ across $\mathcal{S}$ (i.e. $\mathcal{S}\times\mathcal{P}\rightarrow\mathcal{X}$). A logical action $a\in\mathcal{A}$ consists of a precondition $\mathcal{P}_{pre}$ $\langle p_{1},p_{2},\dots\rangle$ and an action’s effect $\mathcal{P}_{eff}=\langle p_{1}^{\prime},p_{2}^{\prime},\dots\rangle$. The precondition represents a set of predicates that must be satisfied for the action to be executed, while the action’s effect describes the change of the resultant state upon action completion. Logical actions define the logical state transitions $\mathcal{X}^{(t)}\times a^{(t)}\rightarrow\mathcal{X}^{(t+1)}$. Thus, any planning problem $\langle\mathcal{S}^{(init)},\mathcal{S}^{(final)}\rangle$ can be formulated as a logical planning problem $\mathcal{Q}=\langle\mathcal{O},\mathcal{D},\mathcal{X}^{(init)},\mathcal{X}^{(final)}\rangle$, which is solved by a symbolic planner to produce a task plan ${a^{(0)},a^{(1)},\dots,a^{(T-1)}}=\texttt{PDDLSolver}(\mathcal{Q})$. Each logical action $a$ must then be integrated with corresponding motion planning skills to generate continuous robotic actions $\tilde{a}$ for execution.
@@ 5 Methodology
Figure 1: Overview of the proposed framework. (1) Human demonstrations, in the form of manipulation trajectories, and the corresponding task descriptions, serve as input. Implementation details is shown in Section˜B.12. (2) PDDLLM initiates thousands of parallel simulations, using the resulting roll-outs and rich physics-based feedback to guide the LLM in summarizing them into meaningful predicates, and returns a predicate library annotated with each predicate’s relevance to the current task. (3) Actions are invented by an LLM that summarizes logical state transition patterns from the demonstration, which is grounded into logical states using the imagined predicates. (4) The predicates and actions are compiled into a planning domain, which is automatically interfaced with motion planning algorithm by the Logical Constrain Adapter (LoCA) to solve new tasks.
@@ 5.1 PDDL Domain Generation
Given a human demonstration and its task description, our system combines simulated physical interactions with LLM-based reasoning to generate an executable PDDL planning domain through predicate imagination and action invention. Simulation verifies the physical feasibility that LLMs alone cannot reliably enforce, while the LLM abstracts these grounded interactions into logical predicates and actions. This combination enables a robust pipeline for planning-domain construction.
@@ 5.1.1 Predicate Imagination
However, the full continuous feature space is highly scattered, making it difficult for the LLM to detect consistent patterns or infer meaningful numeric thresholds. Thus, we discretize the feature space into coherent sub-regions that group similar states together, enabling the LLM to extract stable relational patterns. Additionally, these sub-feature-space boundaries provide physically grounded constraints that allow the motion planner to reliably evaluate predicate truth values during planning.
@@
Figure 2: a. This example illustrates the imagination of predicates for relative object positions. Let $u$ be a configurable variable for each dimension. Object poses are sampled and simulated, with infeasible cases filtered out by the simulation feedback. Feasible subspaces are provided to the LLM to generate first-order predicates with their corresponding physical constraints. Higher-order predicates can be further derived using logical operators (e.g., "not", "for all") from first-order predicates. Diverse predicate examples are provided in Section˜B.9 b. This example shows how the Stack action is invented. Continuous states are grounded into logical states using the imagined predicates, where the state transition represents the logical action. By prompting the LLM with the pair of the current state and the next state after the transition, we obtain the PDDL definition of the action Stack. c. The integration of actions with the motion planner is handled automatically by LoCA, which retrieves the physical constraints associated with each first-order predicate in the action effect set $\mathcal{P}_{eff}$ and applies these constraints for motion planning.
Stage 2. Although the first-order predicates already capture all necessary features of the environment, higher-order predicates are essential for representing more complex relations and improving planning efficiency. Thus, we systematically derive higher-order predicates by combining first-order predicates with logical operators and quantifiers. We focus on using negation operator, together with the quantifiers “for all” ($\forall$) and “there exists” ($\exists$), as they are empirically proven effective in robotics by prior work (Curtis et al., 2022; Silver et al., 2023). These components are combined in all possible valid ways to form richer logical expressions that provide stronger guidance during planning. For example, (is_on ?o1 ?o2) can be negated to produce (not_is_on ?o1 ?o2). When combined with the universal quantifier, this further yields ($\mathrm{}{\forall}$_o1_not_is_on ?o1 ?o2), meaning that for any o1 in the environment, (is_on ?o1 ?o2) is false. This indicates that object o2 is on top.
@@ 5.1.2 Action Invention
5.1.3 Parallel Prompting with Feedback
To avoid planning domain generation failure due to the random nature of LLMs, we adopt a parallel prompting strategy with domain execution feedback. Specifically, multiple candidate PDDL domains are generated from the same demonstration via prompting the LLM multiple times. If a generated domain encounters run-time error (syntax, incomplete domain, fail to reach goals…), then this planning domain is eliminated from the candidates. If more than one planning domain is successfully produced, then a majority vote would be initiated using the LLM to choose the best one. Empirically, we find that when the number of parallel prompts exceeds five, the correctness of the final domain converges to a stable level. Examples of the parallel prompting procedure are shown in Section˜B.7.
5.1.4 Hyperparameter Selection
For each continuous feature $f$, we initialize its discretization scale $u_{f}$ using the smallest non-zero observed difference $d_{min}$ of that feature among all relevant objects. This provides a conservative upper bound on the required $u_{f}$: it ensures that the learned predicates remain expressive enough to capture distinctions for all objects. During LLM-based predicate summarization, the model is allowed to refine the initial discretization hyperparameter $u_{f}$ by either merging or subdividing sub-feature intervals. Specifically, the LLM can merge intervals that are semantically uninformative for the task into a single broader region, and subdivide intervals in feature regions where finer distinctions are necessary. Consequently, even if the initial choice of $u_{f}$ is imperfect, the refinement process enables the LLM to recover the task-relevant granularity and generate predicates that remain both expressive and compact.
@@ 6 Experiments and Baselines
Building on evaluation practices in TAMP, we consider tasks with concrete conditions such as object position, orientation, size, and color (Silver et al., 2023; Garrett et al., 2020; Kumar et al., 2023; Liang et al., 2024; Huang et al., 2025). We structure our experiments along three dimensions, Task Diversity, Task Complexity, and Knowledge Transferability, whose definitions are provided in Section˜B.10.1, with task details described in Section˜B.4. Our experiments are conducted in PyBullet Simulation (Coumans and Bai, 2016), with the symbolic solver from PDDLStream (Garrett et al., 2020) and the motion planning algorithm from PyBullet-Planning (Garrett et al., 2015; Garrett, 2018). We use GPT-4o (OpenAI, 2023), but our method is not tied to this backbone. All experiments use 10 parallel prompts and set $u_{f}=d_{min}$; full ablations are provided in Section˜B.15 and Section˜B.16.
By default, PDDLLM learns a planning domain from a single corresponding demonstration per task. To assess its ability to integrate cross-task knowledge, we also evaluated settings where only demonstrations from simpler related tasks were provided for a more complex one. Specifically, for the rearrangement, the model received only demonstrations of both stacking and unstacking, while for the bridge building, it was given only demonstrations of both stacking and alignment.
@@ 6.1 Baselines and Ablations
We implement six baselines and one ablation of PDDLLM to comprehensively evaluate our method. GPT-4o is used as the default LLM unless otherwise specified, and the motion planning algorithm is kept the same for all methods. The complete details are provided in Section˜B.2
  • LLMTAMP, o1-TAMP, R1-TAMP: LLM-based task and motion planning (LLMTAMP) inspired by Huang et al. (2022a); Li et al. (2022), which use LLMs for task planning, with language description of planning task as input. The task execution demonstration, same as those used in PDDLLM, was provided in the form of natural language in prompt. In addition to GPT-4o, reasoning LLMs, OpenAI’s o1 (OpenAI, 2024) and Deepseek’s R1 (DeepSeek-AI, 2025), are used as backbones for o1-TAMP and R1-TAMP, respectively.
  • LLMTAMP-$F$: Following the method by Huang et al. (2022b); Chen et al. (2024a), LLMTAMP-$F$ extends LLMTAMP with a failure feedback loop.
  • LLMTAMP-FR: Following Wang et al. (2024), LLMTAMP-FR extends failure detection by providing specific failure reasons to guide replanning with the LLM.
@@
Robot planning is required to be real-time in robot deployment, imposing constraints on the planning time allowance (Gammell et al., 2015; Garrett et al., 2015; 2020). In our experiment, a uniform planning time limit of 50 seconds is applied to all planning problems and methods. We measure performance using the planning success rate, as in other studies (Huang et al., 2025; Silver et al., 2023; Kumar et al., 2023). The planning time and token cost are used to measure the planning cost (Zhong et al., 2024). Three parallel runs were conducted to compute the mean and standard error for the planning success rate.